urple Company has $200,000 in net income for 2020 before deducting any compensation or other payment to its sole owner, Kirsten. Kirsten is single and she claims the $12,400 standard deduction for 2020. Purple Company is Kirsten's only source of income. Ignoring any employment tax considerations, compute Kirsten's after-tax income for each of the following situations. Click here to access the 2020 individual tax rate schedule to use for this problem. Assume the corporate tax rate is 21%. When required, carryout intermediate tax computations to the nearest cent and then round your final tax liability to the nearest dollar. a. Purple Company is a proprietorship and Kirsten withdraws $50,000 from the business during the year; Kirsten claims a $37,520 deduction for qualified business income. Kirsten's taxable income is $, and her after-tax income is $. b. Purple Company is a C corporation and the corporation pays out all of its after-tax income as a dividend to Kirsten. Note: Individual taxpayers received preferential treatment regarding the taxation of qualified dividends (0%,15%,20%). For single taxpayers, the 0 percent rate applies to the first $40,000 of taxable income. Purple Corporation's after-tax income is $ and Kirsten's after tax income is $. c. Purple Company is a C corporation and the corporation pays Kirsten a salary of $158,000. Purple Corporation's after-tax income is $ and Kirsten's after-tax income is $.


Question


Asked Sep 16, 2020

272 views

Purple Company has $200,000 in net income for 2020 before deducting any compensation or other payment to its sole owner, Kirsten. Kirsten is single and she claims the $12,400 standard deduction for 2020. Purple Company is Kirsten's only source of income.
Ignoring any employment tax considerations, compute Kirsten's after-tax income for each of the following situations.
Click here to access the 2020 individual tax rate schedule to use for this problem. Assume the corporate tax rate is 21%.
When required, carryout intermediate tax computations to the nearest cent and then round your final tax liability to the nearest dollar.

a. Purple Company is a proprietorship and Kirsten withdraws $50,000 from the business during the year; Kirsten claims a $37,520 deduction for qualified business income.
Kirsten's taxable income is $......., and her after-tax income is $.......

b. Purple Company is a C corporation and the corporation pays out all of its after-tax income as a dividend to Kirsten.
Note: Individual taxpayers received preferential treatment regarding the taxation of qualified dividends (0%,15%,20%). For single taxpayers, the 0 percent rate applies to the first $40,000 of taxable income.
Purple Corporation's after-tax income is $ .....and Kirsten's after tax income is $......

c. Purple Company is a C corporation and the corporation pays Kirsten a salary of $158,000.
Purple Corporation's after-tax income is $ .....and Kirsten's after-tax income is
Answer #1

a)

Kirsten's taxable income is

Income for the year = $200000

Less:

Deduction for qualified business income = $37520

Standard deduction = $12400

Taxable Income = $150,080

Tax on 150,080 = 29,875,70 (14382.50 + (24%*(150080-85525)))

b)

Purple Corporation's after-tax income is

$ 200000-(200000*21%) = $158000

Kristen's taxable income = 158000-12,400 (standard deduction) = $145,600

Kirsten’s tax on $145600 at rates applicable to dividends = (40000*0%)+((145600-40000)*15%) = 15,840

Tax amount = (158000-15,840) = $142,160

c)

Taxable income of corporation = 200000-158000 = 42000

Taxable income of Kristen = 158000-12400 standard deduction = 145600

Kristen’s tax = (14382.50 + (24%*(145600-85525)))= 28,800.50

Kirsten’s after-tax income = 145600-28800.50 = $116,800 (approx)